The diameter of a certain spherical metal ball is 8.23 cm and has a mass of 2.35 kg. It has an enclosed air bubble inside whose diameter is 1.80 cm. What is the metal most likely be made out of?

1
Expert's answer
2021-02-22T10:24:23-0500

Given,

Diameter of ball, D=8.23cm

Radius "R=\dfrac{D}{2}=\dfrac{8.23}{2}=4.115cm"


Radius of Air bubble "r=\dfrac{1.80}{2}=0.9cm"


Mass of ball "m=2.35kg=2350g"

Volume of metal, V = "\dfrac{4}{3}\pi(R^3-r^3)"


"= \dfrac{4}{3}\pi(4.115^3-0.9^3)"


"= 288.82cm^3"

We know,

"Density=\dfrac{Mass}{Volume}"



= "\dfrac{2350}{288.82} g/cm^3"

= "8.136 g\/cm^3"

Hence, the metal is mostly likely to be Copper.